### v1

@@ 5 Concluding remarks
$\overline{\Phi}_{1}^{k-1}(0,1)\leq c^{2}(k+M-1)^{-1}(3M^{-1}+2e^{-1})\leq 4c^{2}k^{-1}.$
Now, we derive the estimates (A.11) and (A.9) by splitting the summations into two parts. Let $\overline{k}_{M}=(k+M-1)/2$. For the estimate (A.11), with the inequality (A.13), there holds
$\Phi_{1}^{k-1}(0,\tfrac{1}{2})=c\sum_{j=1}^{k-1}(k-1-j)^{-\frac{1}{2}}(j+M)^{-1}=c\sum_{j=M+1}^{k+M-1}(k+M-1-j)^{-\frac{1}{2}}j^{-1}\leq c({\rm I}_{11}+{\rm I}_{12}),$
$\mbox{with }{\rm I}_{11}=\sum_{j=M+1}^{[\overline{k}_{M}]}\overline{k}_{M}^{-\frac{1}{2}}j^{-1}\quad\mbox{and}\quad{\rm I}_{12}=\sum_{j=[\overline{k}_{M}]+1}^{2\overline{k}_{M}}(k+M-1-j)^{-\frac{1}{2}}\overline{k}_{M}^{-1}.$
The decomposition is well-defined with the convention $\sum_{j=i}^{i^{\prime}}R_{j}=0$ for any $\{R_{j}\}_{j}$ and $i^{\prime}<i$. Then we have ${\rm I}_{11}\leq\overline{k}_{M}^{-\frac{1}{2}}\ln\overline{k}_{M}\leq\sqrt{2}k^{-\frac{1}{2}}\ln k$ and ${\rm I}_{12}\leq 2\overline{k}_{M}^{-\frac{1}{2}}\leq 2\sqrt{2}k^{-\frac{1}{2}}$. Similarly, for the estimate (A.9), when $\epsilon\in(0,\frac{1}{2}]$, we split $\Phi_{1}^{k_{\rm c}}(M+1,\epsilon+1)$ into
$\Phi_{1}^{k_{\rm c}}(M+1,\epsilon+1)=c\sum_{j=M+1}^{k-2}\big(k+M-1-j\big)^{-(1+\epsilon)}j^{-1}\leq c({\rm I}_{21}+{\rm I}_{22}),$
$\mbox{with }{\rm I}_{21}=\sum_{j=M+1}^{[\overline{k}_{M}]}\overline{k}_{M}^{-(1+\epsilon)}j^{-1}\quad\mbox{and}\quad{\rm I}_{22}=\sum_{j=[\overline{k}_{M}]+1}^{k-2}\big(k+M-1-j\big)^{-(1+\epsilon)}\overline{k}_{M}^{-1}.$
@@
${\rm I}_{21}$ $\leq\overline{k}_{M}^{-(1+\epsilon)}\ln\overline{k}_{M}\leq 2\Big((e\epsilon)^{-1}+2^{\epsilon}(K+1)^{-\epsilon}\ln M\Big)M^{-\epsilon}(k+M-1)^{-1},$
${\rm I}_{22}$ $\leq 2\epsilon^{-1}M^{-\epsilon}(k+M-1)^{-1}.$
Finally, the inequality $(k+M-1)^{-1}\leq c_{K,1}(k+M)^{-1}$ completes the proof of the lemma. ∎
@@
If $c_{0}<C_{0}$ and $K\geq K_{0}$ with sufficiently large $K_{0}$, then $\inf_{\epsilon\in(0,1/2]}f(\epsilon)\leq\frac{3\sqrt{e}}{5}$.
@@
By the definition of $f(\epsilon)$, we have
@@
$\leq$ $6\big(c_{0}\sqrt{L}M\|A\|\epsilon^{-1}(2n\sqrt{L}\|A\|^{-1})^{\frac{\epsilon}{1-\epsilon}}\big)^{1-\epsilon},$
for any $\epsilon\in(0,\frac{1}{2}]$ and $K\geq K_{0}$ with sufficiently large $K_{0}$. Let $g(\epsilon)=\epsilon^{-1}(2n\sqrt{L}\|A\|^{-1})^{\frac{\epsilon}{1-\epsilon}}$. Then
@@
$g(\epsilon)$ attains its minimum over the interval $(0,\frac{1}{2}]$ at $\epsilon=\epsilon^{*}=\frac{2}{c+\sqrt{c^{2}-4}}<\frac{1}{2}$, and $g(\epsilon^{*})=\frac{c+\sqrt{c^{2}-4}}{2}e^{\frac{2(c-2)}{c+\sqrt{c^{2}-4}-2}}\leq ce.$ Thus, for $c_{0}<C_{0}$, we have
$f(\epsilon^{*})\leq 6\big(c_{0}\sqrt{L}M\|A\|ce\big)^{1-\frac{2}{c+\sqrt{c^{2}-4}}}\leq 6\Big(ec_{0}\sqrt{L}M\|A\|\ln(2e^{2}n\sqrt{L}\|A\|^{-1})\Big)^{\frac{1}{2}}<\tfrac{3\sqrt{e}}{5}.$
@@
For the estimate (3.1), if $k\leq K_{0}M$ with some $K_{0}\geq 1$, it holds for any sufficiently large $c_{1}$ and $c_{2}$. Now assume that it holds up to k=KM+t1k=KM+t-1 with some $K\geq K_{0}$ and $1\leq t\leq M-1$. Then we prove the assertion for the case $k=KM+t$. (It holds trivially when $t=0$, since $\Delta_{KM}^{\delta}=0$.) Fix $k=KM+t$ and let $k_{\rm c}:=k-M-2$ and $k_{M}:=[k/M]M=KM$. By the bias-variance decomposition, and the estimates (A.1) and (A.2) in Lemma A.3, we have
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq 2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}k^{-2}+n^{-1}c_{0}^{2}L\|A\|^{2}\Big(M^{2}\overline{\Phi}_{1}^{k_{\rm c}}(M+1,2)+\overline{\Phi}_{k_{\rm c}+1}^{k-1}(0,0)\Big).$
@@
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $c_{K,1}^{2}\Big[5c_{0}^{2}LMn^{-1}\|A\|^{2}(c_{1}+c_{2}\delta^{2})+2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}\Big](k+M)^{-2}$
@@
for any $c_{0}<\|A\|^{-1}(5Ln^{-1}M)^{-\frac{1}{2}}$ and $K\geq K_{0}$, with sufficiently large $K_{0}$ and $c_{1},c_{2}$. Alternatively, using the second estimate in (A.2), we can bound $\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]$ by
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq 2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}k^{-2}+\tfrac{c_{0}L}{2}\Big(8M^{2}\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)+\overline{\Phi}_{k_{\rm c}+1}^{k_{M}-1}(0,1)+\overline{\Phi}_{k_{M}+1}^{k-1}(0,1)\Big).$
Then, with the estimates (A.6) and (A.8), we derive
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $c_{K,1}^{2}\big[c_{0}L(\tfrac{3}{2}+\ln M+4c_{K,2})(c_{1}+c_{2}\delta^{2})+2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}\big](k+M)^{-2}$
@@
for any $c_{0}<L^{-1}(10+\ln M)^{-1}$ and $K\geq K_{0}$, with sufficiently large $K_{0}$ and $c_{1},c_{2}$. This completes the proof of the estimate (3.1).
Next, we prove the estimate (3.2). Similarly, for the cases $k\leq K_{0}M$ with some $K_{0}\geq 1$, the estimate holds trivially for sufficiently large $c_{1}$ and $c_{2}$. Now, assume that the bound holds up to k=KM+t1k=KM+t-1 with some $K\geq K_{0}$ and $1\leq t\leq M-1$, and prove the assertion for the case $k=KM+t$. Fix $k=KM+t$ and let $k_{\rm c}:=k-M-2$ and $k_{M}:=[k/M]M=KM$. By the triangle inequality $\|A\Delta_{k}^{\delta}\|\leq\|\mathbb{E}[A\Delta_{k}^{\delta}]\|+\|\Delta_{k}^{\delta}-\mathbb{E}[A\Delta_{k}^{\delta}]\|,$ and (A.1) and (A.3), we have
@@
with ${\rm I}_{1}=2^{1+\epsilon}\epsilon^{\epsilon}c_{0}^{1-\epsilon}n^{\epsilon}\sqrt{L}\|A\|^{1-2\epsilon}M\Phi_{1}^{k_{\rm c}}(M+1,1+\epsilon).$ By (A.9) (with $c=c_{1}+c_{2}\delta$), we derive
${\rm I}_{1}\leq$ $2^{2+\epsilon}\epsilon^{\epsilon}(c_{1}+c_{2}\delta)c_{0}^{1-\epsilon}n^{\epsilon}\sqrt{L}\|A\|^{1-2\epsilon}M^{1-\epsilon}c_{K,1}c_{K,\epsilon}(k+M)^{-1}=(c_{1}+c_{2}\delta)f(\epsilon)c_{K,1}(k+M)^{-1},$
@@
Then by Lemma A.5 and the inequality $c_{0}\sqrt{L}M\|A\|<200^{-1}$ when $c_{0}<C_{0}$, we derive from (A.16) that
@@
for any $K\geq K_{0}$, with sufficiently large $K_{0}$ and $c_{1},c_{2}$, completing the proof of the theorem. ∎

### v2

@@ Acknowledgments
$\overline{\Phi}_{1}^{\overline{k}}(0,1)\leq c^{2}(\overline{k}+M)^{-1}\big((2+2c_{0}\|B\|)M^{-1}+2e^{-1}\big)\leq(3+2c_{0}\|B\|)c^{2}k^{-1}.$
Now, we derive the estimates (A.11) and (A.9) by splitting the summations into two parts. Let $\widetilde{k_{M}}=(k+M-1)/2$. For the estimate (A.11), there holds
$\Phi_{1}^{\overline{k}}(0,\tfrac{1}{2})\leq$ $\;c\Big(\sqrt{2c_{0}\|B\|}(\overline{k}+M)^{-1}+\sum_{j=1}^{k-2}(\overline{k}-j)^{-\frac{1}{2}}(j+M)^{-1}\Big)$
$=$ $c\Big(\sqrt{2c_{0}\|B\|}(\overline{k}+M)^{-1}+\sum_{j=M+1}^{k+M-2}(\overline{k}+M-j)^{-\frac{1}{2}}j^{-1}\Big)$
$\leq$ $c({\rm I}_{11}+{\rm I}_{12}),$
with
${\rm I}_{11}=\sum_{j=M+1}^{[\widetilde{k_{M}}]}\widetilde{k_{M}}^{-\frac{1}{2}}j^{-1}\quad\mbox{and}\quad{\rm I}_{12}=\bigg(\sqrt{2^{-1}c_{0}\|B\|}+\sum_{j=[\widetilde{k_{M}}]+1}^{2\widetilde{k_{M}}-1}(\overline{k}+M-j)^{-\frac{1}{2}}\bigg)\widetilde{k_{M}}^{-1}.$
The decomposition is well-defined with the convention $\sum_{j=i}^{i^{\prime}}R_{j}=0$ for any $\{R_{j}\}_{j}$ and $i^{\prime}<i$. Then we have
${\rm I}_{11}\leq\widetilde{k_{M}}^{-\frac{1}{2}}\ln\widetilde{k_{M}}\leq\sqrt{2}k^{-\frac{1}{2}}\ln k\quad\mbox{and}\quad{\rm I}_{12}\leq 2\widetilde{k_{M}}^{-\frac{1}{2}}\leq 2\sqrt{2}k^{-\frac{1}{2}}.$
Similarly, for the estimate (A.9), when $\epsilon\in(0,\frac{1}{2}]$, we split $\Phi_{1}^{k_{\rm c}}(M+1,\epsilon+1)$ into
$\Phi_{1}^{k_{\rm c}}(M+1,\epsilon+1)\leq c\sum_{j=M+1}^{k-2}\big(\overline{k}+M-j\big)^{-(1+\epsilon)}j^{-1}\leq c({\rm I}_{21}+{\rm I}_{22}),$
with
${\rm I}_{21}=\sum_{j=M+1}^{[\widetilde{k_{M}}]}\widetilde{k_{M}}^{-(1+\epsilon)}j^{-1}\quad\mbox{and}\quad{\rm I}_{22}=\sum_{j=[\widetilde{k_{M}}]+1}^{k-2}\big(\overline{k}+M-j\big)^{-(1+\epsilon)}\widetilde{k_{M}}^{-1}.$
@@
${\rm I}_{21}$ $\leq\widetilde{k_{M}}^{-(1+\epsilon)}\ln\widetilde{k_{M}}\leq 2\Big((e\epsilon)^{-1}+2^{\epsilon}(K+1)^{-\epsilon}\ln M\Big)M^{-\epsilon}(\overline{k}+M)^{-1},$
${\rm I}_{22}$ $\leq 2\epsilon^{-1}M^{-\epsilon}(\overline{k}+M)^{-1}.$
Finally, the inequality $(\overline{k}+M)^{-1}\leq c_{K,1}(k+M)^{-1}$ completes the proof of the lemma. ∎
@@
If $c_{0}<C_{0}$ and $K\geq M-1$, then $\inf_{\epsilon\in(0,\frac{1}{2}]}f(\epsilon)\leq\frac{4\sqrt{e}}{7}$.
@@
By the definition of $f(\epsilon)$ and the inequality (A.13), we have
@@
$\leq$ $8\big(c_{0}\sqrt{L}M\|A\|\epsilon^{-1}(2n\sqrt{L}\|A\|^{-1})^{\frac{\epsilon}{1-\epsilon}}\big)^{1-\epsilon},$
for any $\epsilon\in(0,\frac{1}{2}]$ and $K\geq M-1$. Let $g(\epsilon)=\epsilon^{-1}(2n\sqrt{L}\|A\|^{-1})^{\frac{\epsilon}{1-\epsilon}}$. Then
@@
$g(\epsilon)$ attains its minimum over the interval $(0,\frac{1}{2}]$ at $\epsilon=\epsilon^{*}=\frac{2}{c+\sqrt{c^{2}-4}}<\frac{1}{2}$, and
$g(\epsilon^{*})=\frac{c+\sqrt{c^{2}-4}}{2}e^{\frac{2(c-2)}{c+\sqrt{c^{2}-4}-2}}\leq ce.$
Thus, for $c_{0}<C_{0}$, we have
$f(\epsilon^{*})\leq$ $8\big(c_{0}\sqrt{L}M\|A\|ce\big)^{1-\frac{2}{c+\sqrt{c^{2}-4}}}$
$\leq$ $8\Big(ec_{0}\sqrt{L}M\|A\|\ln(2e^{2}n\sqrt{L}\|A\|^{-1})\Big)^{\frac{1}{2}}<\tfrac{4\sqrt{e}}{7}.$
@@
If $k\leq K_{0}M$ with some $K_{0}\geq 1$, the estimate (3.1) holds for any sufficiently large $c_{1}$ and $c_{2}$. Now assume that it holds up to k=KM+t1k=KM+t-1 with some $K\geq K_{0}$ and $1\leq t\leq M-1$. Then we prove the assertion for the case $k=KM+t$. (It holds trivially when $t=0$, since $\Delta_{KM}^{\delta}=0$.) Fix $k=KM+t$ and let $k_{\rm c}:=k-M-2$ and $k_{M}:=[\frac{k}{M}]M=KM$. By the bias-variance decomposition, and the estimates (A.1) and (A.2) in Lemma A.3, we have
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}k^{-2}$
$+n^{-1}c_{0}^{2}L\|A\|^{2}\Big(M^{2}\overline{\Phi}_{1}^{k_{\rm c}}(M+1,2)+\overline{\Phi}_{k_{\rm c}+1}^{k-1}(0,0)\Big).$
@@
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $c_{K,1}^{2}\Big[5c_{0}^{2}LMn^{-1}\|A\|^{2}(c_{1}+c_{2}\delta^{2})$
$+2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}\Big](k+M)^{-2}$
@@
for any $c_{0}<3^{-1}\|A\|^{-1}(5Ln^{-1}M)^{-\frac{1}{2}}$ and $K\geq 1$, with sufficiently large $c_{1}$ and $c_{2}$. Alternatively, using the second estimate in (A.2), we can bound $\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]$ by
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}k^{-2}$
$+\tfrac{c_{0}L}{2}\Big(8M^{2}\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)+\overline{\Phi}_{k_{\rm c}+1}^{k_{M}-1}(0,1)+\overline{\Phi}_{k_{M}+1}^{k-1}(0,1)\Big).$
Then, using the estimates (A.6) and (A.8), we derive
$\mathbb{E}[\|A\Delta_{k}^{\delta}\|^{2}]\leq$ $c_{K,1}^{2}\big[c_{0}L(\tfrac{1}{2}+\ln M+2c_{0}\|B\|+4c_{K,2})(c_{1}+c_{2}\delta^{2})$
$+2(\|A\|^{2}\|e_{0}^{\delta}\|^{2}+\delta^{2})M^{2}\big](k+M)^{-2}$
@@
for any $c_{0}<9^{-1}L^{-1}(15+7\ln M)^{-1}$ and $K\geq 1$, with sufficiently large $c_{1}$ and $c_{2}$. This completes the proof of the estimate (3.1).
Next, we prove the estimate (3.2). Similarly, for the cases $k\leq K_{0}M$ with some $K_{0}\geq 1$, the estimate (3.2) holds trivially for sufficiently large $c_{1}$ and $c_{2}$. Now, assume that the bound holds up to k=KM+t1k=KM+t-1 with some $K\geq K_{0}$ and $1\leq t\leq M-1$, and prove the assertion for the case $k=KM+t$. Fix $k=KM+t$ and let $k_{\rm c}:=k-M-2$ and $k_{M}:=[\frac{k}{M}]M=KM$. By the triangle inequality $\|A\Delta_{k}^{\delta}\|\leq\|\mathbb{E}[A\Delta_{k}^{\delta}]\|+\|\Delta_{k}^{\delta}-\mathbb{E}[A\Delta_{k}^{\delta}]\|,$ and (A.1) and (A.3), we have
@@
with
${\rm I}_{1}=2^{1+\epsilon}\epsilon^{\epsilon}c_{0}^{1-\epsilon}n^{\epsilon}\sqrt{L}\|A\|^{1-2\epsilon}M\Phi_{1}^{k_{\rm c}}(M+1,1+\epsilon).$
By (A.9) (with $c=c_{1}+c_{2}\delta$), we derive
${\rm I}_{1}\leq$ $2^{2+\epsilon}\epsilon^{\epsilon}(c_{1}+c_{2}\delta)c_{0}^{1-\epsilon}n^{\epsilon}\sqrt{L}\|A\|^{1-2\epsilon}M^{1-\epsilon}c_{K,1}c_{K,\epsilon}(k+M)^{-1}$
$=$ $(c_{1}+c_{2}\delta)f(\epsilon)c_{K,1}(k+M)^{-1},$
@@
Then by Lemma A.5 and the inequality $c_{0}\sqrt{L}M\|A\|<\big(14^{2}(2+\ln 2)\big)^{-1}$ when $c_{0}<C_{0}$, we derive from (A.16) that
@@
for any $K\geq 35$, with sufficiently large $c_{1}$ and $c_{2}$. This completes the proof of the theorem. ∎